# Secure AI-Driven Super-Resolution for Real-Time Mixed Reality Applications

### v1

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
A significant challenge for 3DoF and 6DoF streaming that does not exist for regular video streaming is motion-to-photon latency, the delay between viewer movement and the updated visuals displayed on a device. High latency can cause motion sickness and reduce immersion in VR and AR[shannigrahi2020next]. Low latency ensures a smoother, more responsive experience. Streaming applications are also bandwidth-intensive since high bandwidth usage may lead to congestion and packet loss, which increase the overall latency.
In our earlier work, we showed that using in-network caching and data reuse can cut down network latency. In this paper, we reduce latency by introducing a novel method that combines super-resolution and Attribute-Based Encryption (ABE) [CPABE]. This approach provides AR/VR streaming applications with the opportunity to trade off between upscaling via super-resolution (SR) [yu2018pu] at the client, data encryption and decryption, and the amount of data transferred over the network.
@@
We evaluate our approach using a point-cloud data set that consists of scenes from offices and living rooms. We use a subset of these scenes to train an SR model that is used at the client for upsampling. In addition, we make use of the ABE-based partial encryption approach we developed for point-cloud streaming [usmani2025securingimmersive360video]. The performance of this system is evaluated based on standard QoE metrics for point-clouds.
Our contribution is that we introduce a novel perspective by integrating security into the machine learning–based super-resolution pipeline for point clouds. Specifically, we combine ABE with point cloud downsampling and ML-based SR. This framework reduces bandwidth usage through downsampling while simultaneously lowering encryption–decryption overhead—an aspect overlooked in prior research. We further develop and train an efficient AI/ML model capable of reconstructing full-resolution point clouds from downsampled data with high fidelity. Experimental results show a nearly linear reduction in network latency, bandwidth consumption, and encryption–decryption time with increased downsampling, while the proposed super-resolution model restores the original resolution with minimal reconstruction error and modest inference time.
@@ II Background
Point clouds represent a collection of discrete points in three-dimensional space, where each point is characterized by its spatial coordinates $(x,y,z)$ and may include additional properties such as color $(r,g,b)$ or surface normals $(n_{x},n_{y},n_{z})$. Combined, these points reconstruct the geometry of an object or scene and are commonly stored in the Polygon File Format (.ply[ply]. Unlike conventional 2D image or video data, point clouds provide a flexible, spatially explicit representation of the physical environment. When captured sequentially over time, these frames form a volumetric video that serves as a core building block for mixed reality (MR) systems that integrate elements of both virtual and augmented environments. Such representations enable 6DoF for immersive user interaction and spatial navigation [10.1145/3394171.3413639], offering a level of realism and interactivity that extends far beyond the three degrees of freedom typical of traditional 360° video experiences [usmani2025securingimmersive360video].
@@ II-B Related Work on Point Cloud Super-Resolution
Super-resolution (SR) using deep learning was first explored for images [Dong2014LearningAD, 8099502] and later extended to conventional 2D videos  [7410425, caballero2017realtimevideosuperresolutionspatiotemporal].The demand for enhancing resolutions eventually led to its application in emerging 3D formats such as point clouds. The pioneering work in learning-based point cloud super-resolution is PU-Net[yu2018pu], which introduced one of the first neural network architectures for point cloud upsampling. Building on this foundation,[10503354] extended SR techniques to LiDAR and large-scale 3D scenes. Other recent studies [10336534] have expanded point cloud SR beyond geometric spatial coordinates to also enhance per-point color attributes, while [10.1007/978-3-030-58529-7_44] introduced a geometrically aware upsampling framework.
@@
There have also been streaming-oriented studies on applying learning based SR to point clouds for volumetric video streaming [10.1145/3446382.3448663, 278306, wang2025volutefficientvolumetricstreaming]. The YuZu framework [278306] represents one of the most comprehensive efforts to integrate SR into volumetric video streaming. Their approach transmits low-resolution point cloud segments and reconstructs higher-resolution versions at the client side using SR models. The system incorporates adaptive bitrate (ABR) control within a DASH-based pipeline and performs evaluations using a client–server setup. Additionally, they introduce a Quality of Experience (QoE) model to assess perceptual quality and demonstrate that SR-based reconstruction achieves significantly better QoE compared to viewport-adaptive methods. More recently, an extension of this work, VoLUT [wang2025volutefficientvolumetricstreaming], introducing a lookup-table (LUT)–based SR approach to further improve efficiency during upscaling.
While prior works have extensively investigated SR-based bandwidth optimization for volumetric and point cloud streaming, our approach introduces a novel perspective by integrating security into this pipeline. Specifically, we combine ABE with point cloud downsampling and ML-based SR. Our framework reduces bandwidth usage through downsampling and lowers encryption–decryption overhead, addressing a dimension that has received little attention in prior work. Moreover, we train an AI/ML model capable of efficiently reconstructing full-resolution point clouds from downsampled data. Although our current focus is on evaluating downsampling, ABE, and reconstruction performance, future work will extend this framework to a full streaming setup with ABR, DASH integration, and QoE evaluation similar to [278306].
@@ II-C Related Work on Attribute-Based Encryption
Secure and flexible access control is a key requirement for collaborative or distributed applications, such as shared 3D or MR data platforms, where users from different organizations or regions access common content. Attribute-Based Encryption (ABE) [CPABE] addresses this need by enabling fine-grained, policy-driven access control directly at the cryptographic layer.
@@
ABE further supports dynamic key and attribute management. If a user’s status changes (for example, Alice leaves the research group), her private key $SK$ can be revoked or allowed to expire without re-encrypting the data. Time-bound revocation can also be enforced by embedding expiration attributes or validity periods in keys, as demonstrated in [reddick2022wip, reddick2021design, reddick2022case]. Administrators can periodically refresh or revoke these attributes to maintain continuous security.
Prior work [10.1007/978-3-319-32689-4_26, 10.1145/3712676.3714450, usmani2025securingimmersive360video] has explored the use of ABE in video streaming, primarily focusing on conventional 2D and 360 video content. The latter studies [10.1145/3712676.3714450, usmani2025securingimmersive360video] demonstrated through real-time streaming evaluations that ABE over HTTP can reduce cache CPU load while maintaining Quality of Experience (QoE) comparable to HTTPS. In particular, [usmani2025securingimmersive360video] extended this concept by integrating selective frame-level encryption with viewport-adaptive protection.
More specifically for point cloud data, [10.1145/3704413.3765298] introduced a selective coordinate encryption mechanism using ABE, enabling encryption of different subsets of spatial coordinates $x,y,z$ to balance computational cost against the desired level of obfuscation. In this work, we build upon and adapt that mechanism to secure point cloud content, as further discussed in Section III-B.
@@ III-B ABE for Securing Point Clouds
To secure the point cloud video streams, we adopt the attribute-based selective coordinate encryption method from our prior work [10.1145/3704413.3765298]. This approach supports encryption at different granularities, encrypting all $X$, $Y$, and $Z$ coordinates, only $X$ and $Y$, or just $X$. The selected coordinates are extracted and encrypted using ABE, with the resulting ciphertext embedded in the point cloud frame as metadata while the original coordinates are removed. During decryption, the metadata is parsed, the encrypted values are recovered, and the original frame is reconstructed by reinserting the decrypted coordinates into their original positions.
In this work, we encrypt all $X$, $Y$, and $Z$ coordinates to evaluate the computational savings achieved through downsampling, while prior work [10.1145/3704413.3765298] analyzed the trade-offs of selective encryption in detail. The use of ABE ensures data confidentiality both in transit and at rest, while avoiding per-client encryption overhead [10.1145/3712676.3714450]. This design also preserves the caching and distribution efficiencies demonstrated in prior streaming studies [10.1145/3712676.3714450, usmani2025securingimmersive360video].
@@ III-C1 Data Preparation
  • For every sparse point $p_{i}=(x_{i},y_{i},z_{i})$, its $K=16$ nearest sparse neighbors are found using a k-dimensional tree (KD-tree). A KD-tree recursively partitions the 3D space along coordinate axes so that each node stores a splitting hyperplane and the points within its subspace. During a query, the tree prunes large portions of the search space by comparing distances to these hyperplanes, thus enabling near–logarithmic-time nearest-neighbor lookups [bentley1975multidimensional]. The mean neighbor distance $\bar{d}_{K}$ obtained from this search acts as a proxy for local sampling density: dense regions yield smaller $\bar{d}_{K}$, while sparse regions yield larger ones.
@@
This process is repeated for all available scenes and environments, producing paired feature–offset samples for each sparse–dense cloud pair. After constructing the dataset of all sparse–dense point correspondences across scenes, we merge them into a unified pool. Each sample retains its semantic tag, e.g., LivingRoom, allowing us to later partition the data into domain-specific subsets for different training configurations. We discuss the dataset composition in Section IV-A.
@@ IV-A Point Cloud Dataset
We utilize two point cloud datasets generated from RGB-D sensor captures, available in Open3D [Open3D]: LivingRoom (56 files) and Office (53 files). Each file is stored in the .ply format and contains spatial coordinates, surface normals, and colors. For simplicity, we remove normals and colors so that each point represents only its spatial position $(x,y,z)$. To ensure strict data separation between training and evaluation, 16 files are excluded from all data preparation and training.
@@
After excluding the sixteen scenes reserved for evaluation, the remaining dense .ply files contain approximately 32.2 million points in total. Each file contains between 0.11–0.89 million points in the LivingRoom domain and 0.11–0.63 million points in the Office domain. From this unified dataset, we randomly sample 5 million pairs for training and 2 million pairs for testing in each configuration. These samples are used to train the Random Forest models described in Sect. III-C2. The dense clouds serve as the ground-truth targets, while their downsampled counterparts provide the sparse inputs used to construct feature–offset pairs.
@@ IV-B Impact of Downsampling on ABE and Data Size
Figure 3(c) illustrates the data sizes (in megabytes) for each point cloud at different resolutions. These sizes include the overhead introduced by ABE. The results demonstrate a linear decrease in size with lower resolutions—for instance, a 49.99% reduction at 50% resolution, 74.99% at 25%, and continuing proportionally for lower levels. This linear reduction in data size also directly translates to a proportional decrease in bandwidth usage during transmission.
@@ IV-C Evaluation of Super-Resolution Model
We evaluate the performance of our AI/ML-based upsampling models by reconstructing high-resolution point clouds from their downsampled counterparts (Sect. IV-A). To quantitatively assess reconstruction fidelity, we employ two standard spatial similarity metrics: the Chamfer Distance (CD) [9879253] and the Hausdorff Distance (HD) [10.1145/3503161.3548384]. The CD measures the average bidirectional nearest-neighbor distance between the reconstructed and reference point clouds, while the HD captures the maximum of these minimum distances, reflecting the worst-case geometric deviation. Lower values of both metrics indicate higher reconstruction fidelity and closer geometric alignment with the original dense point cloud.
@@
Fig. 6 summarizes the per-frame inference times of Models A and B across different input resolutions. The inference runtime of the Random-Forest–based SR model remains lightweight across frames. Direct upsampling from 50% to 100% density completes in under two seconds for most frames, with smaller scenes finishing within one second and the largest (with around 8 M points) taking about three seconds. Multi-stage 2× chaining introduces additional latency because each intermediate stage reconstructs neighborhood structures and re-extracts features for an increasingly dense point set, compounding the cost of KD-tree queries and per-point predictions. An end-to-end model trained to map directly from coarse inputs (e.g., 25% or 12.5%) to full resolution could avoid such redundant feature computations and reduce overall runtime. For mixed-reality pipelines, frame latencies below roughly 20-50 ms are typically considered real-time. Bridging this gap, particularly on resource-constrained clients, would require architectural optimizations such as reusing neighborhood structures across consecutive frames and offloading feature computation to nearby edge servers [alshahrani2020efficient, zhang2022end]. We plan to investigate system-level optimizations for our SR framework in future work.
@@ IV-D Evaluation of Streaming Latency with Downsampling
To evaluate streaming latency, we conducted a simulation on CloudLab [duplyakin2019cloudlab] using two physical nodes: a server hosting pre-encrypted point cloud frames via Apache HTTP and a client running Apache Benchmark (ab) to generate load. The experiment measured the effect of frame resolution and decryption overhead on volumetric streaming latency. We aimed to emulate a realistic streaming setup where 10 clients watch a 30 FPS volumetric video for 600 seconds, resulting in a total of 180,000 HTTP requests (10 x 30 × 600). This workload was issued with concurrency level set to 10, effectively simulating 10 clients fetching frames in parallel.
As the volumetric content, we used a point cloud the dataset (Table I office-28 with 403k points). To model a full-length video, the same frame was repeatedly requested imitating streaming of a point cloud sequence. We conducted four experiments using different resolutions of the point cloud: 100% , 50%, 25%, and 12.5%, obtained by downsampling the original point cloud. Each of these versions was pre-encrypted using ABE, such that the network payload included encryption metadata, ensuring a realistic evaluation of actual streaming conditions with security overhead.

### v2

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/ISM66958.2025.00035
@@ I Introduction
A significant challenge for 3DoF and 6DoF streaming that does not exist for regular video streaming is motion-to-photon latency, the delay between viewer movement and the updated visuals displayed on a device. High latency can cause motion sickness and reduce immersion in VR and AR[20]. Low latency ensures a smoother, more responsive experience. Streaming applications are also bandwidth-intensive since high bandwidth usage may lead to congestion and packet loss, which increase the overall latency.
In our earlier work, we showed that using in-network caching and data reuse can cut down network latency [25, 26]. In this paper, we reduce latency by introducing a novel method that combines super-resolution and Attribute-Based Encryption (ABE) [3]. This approach provides AR/VR streaming applications with the opportunity to trade off between upscaling via super-resolution (SR) [28] at the client, data encryption and decryption, and the amount of data transferred over the network.
@@
We evaluate our approach using a point-cloud data set that consists of scenes from offices and living rooms. We use a subset of these scenes to train an SR model that is used at the client for upsampling. In addition, we make use of the ABE-based partial encryption approach we developed for point-cloud streaming [26]. The performance of this system is evaluated based on standard QoE metrics for point-clouds.
Our contribution is that we introduce a novel perspective by integrating security into the ML-based super-resolution pipeline for point clouds. Specifically, we combine ABE with point cloud downsampling and ML-based SR. This framework reduces bandwidth usage through downsampling while simultaneously lowering encryption-decryption overhead, an aspect overlooked in prior research. We further develop and train an efficient AI/ML model capable of reconstructing full-resolution point clouds from downsampled data with high fidelity. Experimental results show a nearly linear reduction in network latency, bandwidth consumption, and encryption-decryption time with increased downsampling, while the proposed super-resolution model restores the original resolution with minimal reconstruction error and modest inference time.
@@ II Background
Point clouds represent a collection of discrete points in three-dimensional space, where each point is characterized by its spatial coordinates $(x,y,z)$ and may include additional properties such as color $(r,g,b)$ or surface normals $(n_{x},n_{y},n_{z})$. Combined, these points reconstruct the geometry of an object or scene and are commonly stored in the Polygon File Format (.ply[23]. Unlike conventional 2D image or video data, point clouds provide a flexible, spatially explicit representation of the physical environment. When captured sequentially over time, these frames form a volumetric video that serves as a core building block for mixed reality (MR) systems that integrate elements of both virtual and augmented environments. Such representations enable 6DoF for immersive user interaction and spatial navigation [22], offering a level of realism and interactivity that extends far beyond the three degrees of freedom typical of traditional 360° video experiences [26].
@@ II-B Related Work on Point Cloud Super-Resolution
Super-resolution (SR) using deep learning was first explored for images [7, 11] and later extended to conventional 2D videos  [12, 5]. The demand for enhancing resolutions eventually led to its application in emerging 3D formats such as point clouds. The pioneering work in learning-based point cloud super-resolution is PU-Net[28], which introduced one of the first neural network architectures for point cloud upsampling. Building on this foundation,[19] extended SR techniques to LiDAR and large-scale 3D scenes. Other recent studies [9] have expanded point cloud SR beyond geometric spatial coordinates to also enhance per-point color attributes, while [15] introduced a geometrically aware upsampling framework.
@@
There have also been streaming-oriented studies on applying learning based SR to point clouds for volumetric video streaming [29, 30, 27]. The YuZu framework [30] represents one of the most comprehensive efforts to integrate SR into volumetric video streaming. Their approach transmits low-resolution point cloud segments and reconstructs higher-resolution versions at the client side using SR models. The system incorporates adaptive bitrate (ABR) control within a DASH-based pipeline and performs evaluations using a client-server setup. Additionally, they introduce a Quality of Experience (QoE) model to assess perceptual quality and demonstrate that SR-based reconstruction achieves significantly better QoE compared to viewport-adaptive methods. More recently, an extension of this work, VoLUT [27], introducing a lookup-table (LUT)-based SR approach to further improve efficiency during upscaling.
While prior works have extensively investigated SR-based bandwidth optimization for volumetric and point cloud streaming, our approach introduces a novel perspective by integrating security into this pipeline. Specifically, we combine ABE with point cloud downsampling and ML-based SR. Our framework reduces bandwidth usage through downsampling and lowers encryption-decryption overhead, addressing a dimension that has received little attention in prior work. Moreover, we train an AI/ML model capable of efficiently reconstructing full-resolution point clouds from downsampled data. Although our current focus is on evaluating downsampling, ABE, and reconstruction performance, future work will extend this framework to a full streaming setup with ABR, DASH integration, and QoE evaluation similar to [30].
@@ II-C Related Work on Attribute-Based Encryption
Secure and flexible access control is a key requirement for collaborative or distributed applications, such as shared 3D or MR data platforms, where users from different organizations or regions access common content. Attribute-Based Encryption (ABE) [3] addresses this need by enabling fine-grained, policy-driven access control directly at the cryptographic layer.
@@
ABE further supports dynamic key and attribute management. If a user’s status changes (for example, Alice leaves the research group), her private key $SK$ can be revoked or allowed to expire without re-encrypting the data. Time-bound revocation can also be enforced by embedding expiration attributes or validity periods in keys, as demonstrated in [18, 16, 17]. Administrators can periodically refresh or revoke these attributes to maintain continuous security.
Prior work [13, 25, 26] has explored the use of ABE in video streaming, primarily focusing on conventional 2D and 360 video content. The latter studies [25, 26] demonstrated through real-time streaming evaluations that ABE over HTTP can reduce cache CPU load while maintaining Quality of Experience (QoE) comparable to HTTPS. In particular, [26] extended this concept by integrating selective frame-level encryption with viewport-adaptive protection.
More specifically for point cloud data, [24] introduced a selective coordinate encryption mechanism using ABE, enabling encryption of different subsets of spatial coordinates $x,y,z$ to balance computational cost against the desired level of obfuscation. In this work, we build upon and adapt that mechanism to secure point cloud content, as further discussed in Sect. III-B.
@@ III-B ABE for Securing Point Clouds
To secure the point cloud video streams, we adopt the attribute-based selective coordinate encryption method from our prior work [24]. This approach supports encryption at different granularities, encrypting all $X$, $Y$, and $Z$ coordinates, only $X$ and $Y$, or just $X$. The selected coordinates are extracted and encrypted using ABE, with the resulting ciphertext embedded in the point cloud frame as metadata while the original coordinates are removed. During decryption, the metadata is parsed, the encrypted values are recovered, and the original frame is reconstructed by reinserting the decrypted coordinates into their original positions.
In this work, we encrypt all $X$, $Y$, and $Z$ coordinates to evaluate the computational savings achieved through downsampling, while prior work [24] analyzed the trade-offs of selective encryption in detail. The use of ABE ensures data confidentiality both in transit and at rest, while avoiding per-client encryption overhead [25]. This design also preserves the caching and distribution efficiencies demonstrated in prior streaming studies [25, 26].
@@ III-C1 Data Preparation
  • For every sparse point $p_{i}=(x_{i},y_{i},z_{i})$, its $K=16$ nearest sparse neighbors are found using a k-dimensional tree (KD-tree). A KD-tree recursively partitions the 3D space along coordinate axes so that each node stores a splitting hyperplane and the points within its subspace. During a query, the tree prunes large portions of the search space by comparing distances to these hyperplanes, thus enabling near–logarithmic-time nearest-neighbor lookups [2]. The mean neighbor distance $\bar{d}_{K}$ obtained from this search acts as a proxy for local sampling density: dense regions yield smaller $\bar{d}_{K}$, while sparse regions yield larger ones.
@@
This process is repeated for all available scenes and environments, producing paired feature-offset samples for each sparse-dense cloud pair. After constructing the dataset of all sparse-dense point correspondences across scenes, we merge them into a unified pool. Each sample retains its semantic tag, e.g., LivingRoom, allowing us to later partition the data into domain-specific subsets for different training configurations. We discuss the dataset composition in Sect. IV-A.
@@ IV-A Point Cloud Dataset
We utilize two point cloud datasets generated from RGB-D sensor captures, available in Open3D [14]: LivingRoom (56 files) and Office (53 files). Each file is stored in the .ply format and contains spatial coordinates, surface normals, and colors. For simplicity, we remove normals and colors so that each point represents only its spatial position $(x,y,z)$. To ensure strict data separation between training and evaluation, 16 files are excluded from all data preparation and training.
@@
After excluding the sixteen scenes reserved for evaluation, the remaining dense .ply files contain approximately 32.2 million points in total. Each file contains between 0.11–0.89 million points in the LivingRoom domain and 0.11–0.63 million points in the Office domain. From this unified dataset, we randomly sample 5 million pairs for training and 2 million pairs for testing in each configuration. These samples are used to train the Random Forest models described in Sect. III-C2. The dense clouds serve as the ground-truth targets, while their downsampled counterparts provide the sparse inputs used to construct feature-offset pairs.
@@ IV-B Impact of Downsampling on ABE and Data Size
Fig. 3(c) illustrates the data sizes (in megabytes) for each point cloud at different resolutions. These sizes include the overhead introduced by ABE. The results demonstrate a linear decrease in size with lower resolutions—for instance, a 49.99% reduction at 50% resolution, 74.99% at 25%, and continuing proportionally for lower levels. This linear reduction in data size also directly translates to a proportional decrease in bandwidth usage during transmission.
@@ IV-C Evaluation of Super-Resolution Model
We evaluate the performance of our AI/ML-based upsampling models by reconstructing high-resolution point clouds from their downsampled counterparts (Sect. IV-A). To quantitatively assess reconstruction fidelity, we employ two standard spatial similarity metrics: the Chamfer Distance (CD) [21] and the Hausdorff Distance (HD) [10]. The CD measures the average bidirectional nearest-neighbor distance between the reconstructed and reference point clouds, while the HD captures the maximum of these minimum distances, reflecting the worst-case geometric deviation. Lower values of both metrics indicate higher reconstruction fidelity and closer geometric alignment with the original dense point cloud.
@@
Fig. 6 summarizes the per-frame inference times of Models A and B across different input resolutions. The inference runtime of the Random-Forest-based SR model remains lightweight across frames. Direct upsampling from 50% to 100% density completes in under two seconds for most frames, with smaller scenes finishing within one second and the largest (with around 800K points) taking about three seconds. Multi-stage 2× chaining introduces additional latency because each intermediate stage reconstructs neighborhood structures and re-extracts features for an increasingly dense point set, compounding the cost of KD-tree queries and per-point predictions. An end-to-end model trained to map directly from coarse inputs (e.g., 25% or 12.5%) to full resolution could avoid such redundant feature computations and reduce overall runtime. For mixed-reality pipelines, frame latencies below roughly 20-50 ms are typically considered real-time. Bridging this gap, particularly on resource-constrained clients, would require architectural optimizations such as reusing neighborhood structures across consecutive frames and offloading feature computation to nearby edge servers [1, 31]. We plan to investigate system-level optimizations for our SR framework in future work.
@@ IV-D Evaluation of Streaming Latency with Downsampling
To evaluate streaming latency, we conducted a simulation on CloudLab [8] using two physical nodes: a server hosting pre-encrypted point cloud frames via Apache HTTP and a client running Apache Benchmark (ab) to generate load. The experiment measured the effect of frame resolution and decryption overhead on volumetric streaming latency. We aimed to emulate a realistic streaming setup where 10 clients watch a 30 FPS volumetric video for 600 seconds, resulting in a total of 180,000 HTTP requests (10 x 30 × 600). This workload was issued with concurrency level set to 10, effectively simulating 10 clients fetching frames in parallel.
As the volumetric content, we used a point cloud the dataset (Table I office-28 with 430K points). To model a full-length video, the same frame was repeatedly requested imitating streaming of a point cloud sequence. We conducted four experiments using different resolutions of the point cloud: 100% , 50%, 25%, and 12.5%, obtained by downsampling the original point cloud. Each of these versions was pre-encrypted using ABE, such that the network payload included encryption metadata, ensuring a realistic evaluation of actual streaming conditions with security overhead.